Circular No. 7328

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gh IN NGC 2986 S. Nakano, Sumoto, Japan, reports the discovery by Kesao Takami-

zawa, Saku-machi, Nagano-ken, of a supernova (mag 14.6) on two T-Max 400-120 films taken with 5-min exposures on Dec. 3.806 UT (and 6 min later) with a 0.10-m f/4.0 lens. Takamizawa provides a position for SN 1999gh of α = 9^{h}44^{m}20^{s}.2, δ = −21^{o}16′13′′ (equinox 2000.0), which is eastnortheast of the center of NGC 2986. Nothing is visible on 14 patrol films (limting mag 14.5–15.6) taken by Takamizawa between 1995 Jan. 1 and 1998 Nov. 26. Nakano also forwards the following precise position end figures, measured by Y. Kushida (Yatsugatake South Base Observatory) from an unfiltered CCD frame taken by R. Kushida with a 0.40-m f/6.3 Schmidt-Cassegrain reflector on Dec. 5.819, when the supernova was at mag 14.8: 19^{s}.75, 25′′.0 (which is 52′′ east and 15′′.8 north of the galaxy’s center). Nothing is visible on previous CCD images taken by Kushida between 1996 Dec. and 1999 Feb. (limiting mag 18).

A. V. Filippenko, University of California at Berkeley, and P. Garnavich, Harvard-Smithsonian Center for Astrophysics, report that a CCD spectrum (range 390–1040 nm) obtained on Dec. 5 UT with the Keck-II telescope shows that SN 1999gh is of type Ia, near maximum brightness. The photospheric velocity, measured from the minimum of the Si II 615-nm absorption, is 11 200 km/s. The ratio of the 580-nm absorption depth to that of the Si II 615-nm feature is similar to that seen in SN 1992A, which was somewhat less luminous than normal. Note that the host galaxy has Hubble type E2.

SUPERNOVAE 1999fz, 1999gb, 1999gd, 1999gf Filippenko and Garnavich also report spectral confirmations of super-

novae, obtained as above. SN 1999fz in UGC 8164 (cf. IAUC 7314) is of type Ia, probably ∼ 1.5–2 months past maximum brightness. SN 1999gb in NGC 2532 (cf. IAUC 7316) is indeed of type IIn, as noted on IAUC 7326. SN 1999gd in NGC 2623 (cf. IAUC 7319) is superposed on a tidal tail in the interacting host galaxy; it is of type Ia, near maximum brightness, and exhibits prominent, narrow Na D absorption, suggesting the possibility of significant extinction. SN 1999gf in UGC 5515 (cf. IAUC 7321) is of type Ia, probably ∼ 2 months past maximum brightness.

1999 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green